# inkscapepath=inkscape, inkscapearea=page, clean=true, tex=false

### v1

@@ 3.2 Image-Level Analysis: Trends and Representations
To gain a deeper understanding of the generated visual material beyond the initial impression of the user-facing visual layer of the CivitAI platform, we examine all the images and their metadata shared on the site between January 2023 and December 2024. In our analysis, we first focus on quantifying temporal trends in the distribution of NSFW classification levels assigned to generated images by the CivitAI platform and extracted from the available image metadata. Secondly, we analyze the images themselves in relation to the demographic characteristics of the depicted subjects. Specifically, we apply a classifier trained to infer age and gender in order to uncover broader trends in the generation of human figures. We focus on age and gender because they are fundamental markers of social identity whose representation in synthetic media can illuminate patterns of discrimination, idealization and exclusion.
@@ 3.2.1 Quantifying NSFW Trends
In the following section, we document how the composition of the CivitAI image corpus changed over time with respect to the corresponding NSFW levels associated with images. A previous study described by Palmini et al. (2024), revealed a disproportionate increase in NSFW imagery compared to SFW on CivitAI. The NSFW ratio, defined as the proportion of all images flagged as NSFW and assigned any browsing level other than 1, which is SFW, increased from 56 % in October 2023 to 73 % in April 2024. Additionally, Wei et al. (2024) demonstrated that NSFW images receive more social engagement on CivitAI than SFW images. Hence, we expected this ratio to further increase. Our current study extends the dataset to include model metadata and covers a full two-year period, from January 2023 to December 2024. With this extended dataset, we confirm a continuation of the trend, with the NSFW ratio increasing from 41 % in January 2023 to reaching 80 % by the end of December 2024. This trend is visualized in Figure 4, which presents a stacked bar diagram showing the color-coded proportions of over 40 million images according to their assigned NSFW browsing level, posted to CivitAI between January 2023 and December 2024. A more detailed account of monthly trends in image generation and NSFW classification is provided in Table  1 in the Supplementary Material. This initial analysis reveals that NSFW content is not only growing in absolute volume on CivitAI but is also becoming increasingly dominant within the platform’s overall visual output. The rise in NSFW content can be seen as symptomatic of an algorithmic attention economy, in which content that provokes and sensationalizes is more likely to be amplified due to its capacity to capture user engagement. Despite CivitAI’s implementation of a five-tier content rating system and its combined use of automated detection tools, human moderation, and community reporting, the platform’s NSFW content continues to expand at a rapid pace. To gain deeper insight into the representational patterns within this content, we conduct a face classification analysis focusing on inferred gender and age.
@@
Figure 4: Inferred age and gender predictions using MiVOLO on a 0.1 % sample of temporally representative images from the Civiverse-40M dataset.
@@ 3.2.2 Demographic Patterns in Generated Images
To analyze and quantify the demographic distribution of human subjects depicted in generated images shared via CivitAI, we employed the MiVOLO (Multi-Input VOLO) model introduced by Kuprashevich and Tolstykh (2023). This is a transformer-based system designed to predict age and gender from images, even when faces are partially or fully occluded, while excluding images without human subjects. The results are shown in Figure 4, which shows the distribution of estimated gender and ages and the corresponding NSFW browsing levels. The left panel shows the distribution of data subjects classified as female, and the right panel those classified as male. The x-axis represents the predicted subject age, rounded to the nearest half-year, while the y-axis indicates the number of detected subjects. Each square marker is color-coded based on the NSFW levels color scheme introduced in Figure  4. Markers are stacked vertically to reflect subject density within each age group and highlight the distribution of NSFW levels by color.
@@
The gender and age estimation was performed on a sampled subset of the original Civiverse-40M dataset. The subset, consists of 40,636 images, representing approximately 0.1 % of the full dataset. It was created by sampling one out of every 1000 images across the two-year period, ensuring a temporally balanced and distributionally representative sample. Using the MiVOLO classifier, we identified 29,073 human subjects in 40,636 images, discarding images with no human subjects or a low confidence level for gender classification. The results of gender classification indicate a massive overrepresentation of female subjects, with an average female-to-male ratio of 6.24 to 1 across the dataset. NSFW co-occurrence rates were also notably higher for female subjects, with 85.44 % of images depicting women with a browsing level other than 1, compared to male subjects, where the NSFW incidence was 73.66 at %. The average age for female subjects was 23.67 years (with a standard deviation of 5.1), while male subjects had a higher average age of 31.16 years with a standard deviation of 9.09 years. This indicates a narrower age range and a strong emphasis on sexualized portrayals of younger women, in contrast to a broader age distribution and more varied depictions of male individuals. A more detailed overview of these results, including a monthly breakdown of estimated gender and age distributions in sampled images, proportions across NSFW browsing levels, female-to-male ratios, and mean age statistics for each group is given in Table 2 in the Supplementary Material.
@@
It is important to note several limitations inherent to this analysis. As described in Section 3.1, CivitAI assigns browsing levels using a combination of Amazon Rekognition, unspecified open-source NSFW classifiers, and manual review. NSFW classifiers are known to disproportionately misclassify images depicting female subjects as explicit when compared to male subjects (Leu et al., 2024). This may slightly skew the distribution of NSFW labels across gender groups, particularly by overrepresenting female subjects in higher-sensitivity categories. It is important to acknowledge the limitations of using MiVOLO’s binary gender classification, which relies on visually inferred physical traits and does not account for non-binary or gender-diverse identities. Additionally, commercial classification systems such as MiVOLO have been shown to produce uneven error rates across demographic groups, particularly for individuals with darker skin tones (Buolamwini and Gebru, 2018). Age estimation regression is also subject to known forms of encoded discrimination and biases, e.g. a study by Chu et al. (2022) has shown that age predictors tend to introduce, what has been described as digital ageism, which can lead to the under representation or distortion of older adults, especially older women. Nevertheless, our results provide meaningful insights into the demographic and content-level dynamics of AI-generated imagery on CivitAI. Despite the limitations of current classification systems, the analysis reveals a pronounced overrepresentation of female subjects, particularly in NSFW-labeled content, as well as a consistent emphasis on youthful female appearances.
@@ 3.3 Model-Level Analysis: Themes and Biases
While generated images often receive the most attention, it is important to recognize that these are fundamentally model-sharing platforms, where the distribution and adaptation of models can reinforce ethically concerning trends, including the hypersexualization and objectification of women, often in a more systematic and enduring way than individual images. Therefore, in the following sections, we shift our focus from the images shared on CivitAI to the analysis of generative models involved in their genesis. In our analysis, we distinguish between different types of models shared on CivitAI. Of the 232,164 assets published to CivitAI between November 2022 and January 2025, 4.28 % were standalone models of the type checkpoint, while 94.04 % were model adapters, which can only be used in conjunction with a compatible base model of the type checkpoint, to generate images. A detailed overview of different asset types shared on CivitAI is given in  Figure 12 in the Supplementary Material. To better understand the thematic focus of the models shared on CivitAI, we first analyze promotional tags, which are user-generated labels that describe the intended use of specific models and model adapters. We then focus on examining checkpoints models, conducting an inference experiment on the ten most popular models to evaluate the reproduction of gender biases. Next, we analyze model adapters, such as LoRA models, by looking at the thirty most popular examples on the platform and assessing their role in fine-tuning or extending base models. Within the broader category of model adapters, we also identify a specific subset designed to mimic the visual likeness of a particular individual. We refer to this group as deepfake adapters, which make up roughly 15 % of all model adapters on the platform.
@@ 3.3.1 Thematic Overview of Models shared on CivitAI
To gain insight into the dominant thematic orientations, we collect promotional tags added by the creators to the various types of models shared on CivitAI. Promotional tags are labels that categorize and describe the thematic, aesthetic, and functional attributes of models and adapters shared on the platform, assigned by users during the model creation and publishing process. We analyze the frequency and co-occurrence of these tags across the platform . These frequencies are visualized in Figure 5 (interactive version available [^11] An interactive version of this graph is available here: https://pm-paper-viz.github.io/visualizations/figure_5.html.), which shows a co-occurrence network of the top 300 most common tags. Each node represents a promotional tag, with node size reflecting its relative frequency of use. Edges between nodes indicate co-occurrence of tags, with thicker and darker lines denoting stronger co-occurrence rates.
To aid visual interpretation, the size of nodes is normalized and the top three most frequent nodes indicate their absolute occurrence counts. For instance, the most prevalent tag, “character” occurs 112,051 times, approximately in half of all the models across the dataset. The “character” tag occurs together with the tag “anime” in 53,792 instances. Anime aesthetics overall seem to be a prevalent theme on the platform with the tag “anime” appearing in roughly 30 % of models. Additionally, tags denoting gender are highly prevalent. The tag “woman” appears in 29.2 % of shared models, “girls” in 15.7 %, and “female” in 11.2 %. Taken together, these figures indicate a pronounced emphasis on women and girls as depicted subjects across all assets exchanged on the platform. Additionally, these tags are commonly associated with tags such as “sexy”, “pose”, certain hairstyles and attire and visual characteristics such as portraiture. The conceptual landscape is dominated by several key thematic clusters. One major cluster focuses on visual style, with tags related to mediums such as oil painting, watercolor, photorealism, sketching, and cinematic composition. The most active cluster revolves around character and anime and is connected to franchises like Dragon Ball, One Piece, and Gundam, and descriptors like TV show, visual novel, and game character. This cluster reflects the strong influence of fan culture and pop media. Among the various thematic patterns observed, the prominence of gendered tags stands out as one of the most striking results which highlights deeper structural issues within the platform’s content ecosystem. Tags such as “woman” and “girl” not only dominate quantitatively but are also frequently linked to sexualized representations, reinforcing long-standing patterns of objectification.
@@ 3.3.2 Bias in Model Checkpoints
To visualize what embedded discrimination in popular diffusion models means in terms of visual output, we performed an experiment on the top ten most popular checkpoints ranked by download count. Checkpoints are saved versions of a pretrained model, containing all necessary parameters to perform inference tasks without additional training and serving as the foundational assets for image synthesis, whereas adapters such as LoRA function as extensions merged into checkpoints to tweak their generative capabilities. Checkpoints currently account for roughly 6 % of shared assets, however as they are the starting point to generate images while adapters are optional, they mostly account for the top assets according to download count. To assess gender bias in popular checkpoint models shared on CivitAI, we performed a simple inference experiment based on comparing outputs of two different neutral text prompts. We selected the ten most popular checkpoint models according to download count available on the platform as of January 19th, 2025. For instance, Realistic Vision, the most downloaded asset at the time, had been downloaded over 1.5 million times, illustrating the scale of model adoption and its potential impact on generated content. We prompted each checkpoint model separately with the single-word prompts “woman” and “man”. Prompt embeddings were generated using the standard CLIP text encoder expected by each checkpoint. All images were generated using consistent parameters: CFG scale 6.5, Euler sampler Qi et al. (2024), and 15 inference steps. To ensure comparability across outputs, we initialized the generation process with the same random noise seed for each model. The selected checkpoints and their inference results are visualized in Figure 6. For comparison, we also included results from Stable Diffusion 1.5 and Stable Diffusion XL 1.0, since most of the selected checkpoints, (aside from Pony Diffusion), are derivatives of these models. The generated outputs across all models revealed a stark difference in gender representation: women were predominantly depicted nude, while men were frequently portrayed in professional attire.
@@
While the baseline models, Stable Diffusion 1.5 and Stable Diffusion XL, still exhibit noticeable gender biases, these biases are even more pronounced in the model derivatives popular on CivitAI. It must be noted that our selection of checkpoint models was based on platform popularity metrics without restricting for model generality or application domain. As a result, the sample includes both general-purpose models and specialized models (e.g., anime-stylized, NSFW-oriented). This choice reflects the actual diversity of models encountered by end-users on CivitAI and allows us to capture platform-wide trends in gender representation. However, it also introduces specific variations that may intensify gender biases. Taken together, these findings point toward embedded discrimination within the platform’s model ecosystem, where user preferences, model fine-tuning practices, and training data biases converge to systematically reproduce and amplify stereotypical and hypersexualized portrayals of women.
@@ 3.3.3 Most Popular Model Adapters
A brief examination of the promotional tags of the most popular models confirms the thematic orientation identified in our broader tag analysis, with frequent references to tags such as “anime”, “female” and “sexy”, suggesting a strong emphasis on gendered and sexualized representations. Indeed, a closer look at the 30 most downloaded adapters reveals a strong presence of adapters implying pornographic themes. Many of the promotional tags accompanying these models specifically emphasize female subjects or body parts, with three models focused solely on the depiction of vaginas. Disconcertingly, one of the most popular adapters, Nudify XL: Better Bodies, directly references the practice of non-consensual “nudification” where photos of dressed individuals are manipulated to appear naked.
Interestingly, some of the model adapters that are not explicitly intended for generating ethically concerning content, still contain questionable tags. For example, the model named Not Artists Styles for Pony Diffusion, was created for the purpose of emulating artistic styles such as oil painting, watercolor, or anime, but it was trained on synthetic data to avoid directly replicating the styles of real, specific artists. Surprisingly, even a model whose primary purpose is to offer stylistic variation includes the tag “female” among its promotional labels. One possible explanation is that the inclusion of this tag may enhance the visibility of the model. While CivitAI does not provide formal documentation on how tags influence model visibility, it is reasonable to conclude that widely used tags may enhance discoverability through the platform’s search and filtering mechanisms. Given that users can browse and search models by tag, including popular terms likely increases the chances of a model being encountered. Therefore, the inclusion of the tag “female” may reflect both the content of sample outputs and a strategic effort to align with common user search behaviors. This can create a reinforcing loop where the repeated use of tags such as “female” “girl” or “sexy” both reflects and amplifies their dominance on the platform. When users frequently tag models with gendered terms, these tags become more visible and searchable, encouraging other creators to include the same tags to gain exposure or align with perceived user interests. Over time, this behavior can solidify certain themes, especially those centered around female representation, as defaults or norms within the platform.
@@ 3.3.4 Deepfake Adapters and Targeted Subjects
One of the most ethically concerning applications of model personalization is the creation of deepfake-related model adapters intended to replicate real individuals. To better understand the scope and impact of this phenomenon, we focused on analyzing the subset of deepfake adapters, investigating their prevalence, common use cases, and the demographic characteristics of the individuals they target. We identified a significant subset of models specifically designed to replicate real individuals, often celebrities, influencers, streamers or public figures whose appearances are accessible online. According to Deeptrace’s 2019 report, 96 % of all deepfake content online was pornographic (Ajder et al., 2019), a trend that intensified by 2023 when a report by Home Security Heroes found that 98 % of deepfake videos were pornographic and 99 % of the victims were women (Home Security Heroes, 2023; Birrer and Just, 2024). Although CivitAI prohibits users from publicly sharing explicit content involving real people, it has no control over how models are used once downloaded or what is shared on external platforms. This creates a significant blind spot in the enforcement of ethical guidelines, as deepfake models can still be deployed in harmful ways beyond the platform’s visibility. Given this regulatory gap and the continuing risk of abuse, we sought to understand who is being targeted by these models. To gain more insight into the individuals currently targeted by the practice of non-consensual deepfakes,we annotated our dataset of deepfake adapters using a large language model, similarly to the approach pioneered by Wei et al. (2024). This allowed us to identify the most commonly associated professions and countries linked to those individuals.
@@
To visualize our results, we use sunburst charts, which represent nested categories as concentric rings, to illustrate how inferred professions are distributed across gender and country associations. Figure 8 includes two sunburst charts that visualize the distribution of inferred professions among individuals represented in the dataset. Figure  8(a) shows this distribution disaggregated by gender, highlighting the prominence of certain professions across male and female categories. While Figure  8(b) breaks down the same professional categories by the most frequently associated countries, offering information on the geographical concentration of targeted individuals.
Our results indicate that actors emerged as the most common profession across both genders and all countries, with the exception of Japan. Among women, adult performers were the second most frequent category, whereas public figures ranked second among men. The majority of targeted individuals were associated with countries in the Global North, specifically the United States, the United Kingdom, and Japan. Cultural factors appear to influence who is targeted: for example, Japan showed a notably high presence of voice actors compared to other countries, likely reflecting the country’s strong anime and gaming industries. Sports professionals were also frequently identified, a pattern that may be attributed to the increased media exposure surrounding international events such as the 2024 Olympics.
Although this analysis gives some insight into the cultural and professional context of the victims of deepfake, it also relies on several simplifying assumptions and methodological constraints. Language models may misclassify individuals or fail to capture multiple roles. We restricted the prediction of professions to three options and selected only the first one for the final categorization,which necessarily simplifies the complexity of real-world identities. Geographic bias is also a potential issue, as DeepSeek may conflate politically or culturally distinct regions. Despite these methodological constraints, our analysis provides a structured entry point for understanding the social dynamics underlying deepfake generation.
The creation of deepfake explicit images of real people is a clear abuse that clearly illustrates the dangers of model personalization. However, beyond these openly malicious uses lies a subtler, systemic issue: the propagation of misogynistic and ethically problematic content through the foundational structures of model training and configuration. These problems do not necessarily stem from explicit intent, but rather from seemingly neutral mechanisms, such as frameworks for automatically associating images with textual descriptions w are used for training models. In the next section, we shift our focus to these less visible yet deeply embedded issues, examining how underlying training mechanisms encode gendered, sexualized, and ethically troubling assumptions into personalized models.

### v2

@@ 2.2 Image-Level Analysis: Trends and Representations
To deepen our understanding of the visual material beyond initial impressions of the user-facing layer of the CivitAI platform, we examined all images and their metadata shared between January 2023 and December 2024. We first quantified temporal trends in the distribution of NSFW classification levels assigned by the CivitAI platform, as extracted from image metadata. We then analyzed the images themselves with regard to the demographic characteristics of depicted subjects, employing a classifier trained to infer age and gender to reveal broader trends in the generation of human figures.
@@ 2.2.1 Quantifying NSFW Trends
A previous study described by Palmini et al. (2024) revealed a disproportionate increase in NSFW imagery compared to SFW on CivitAI. The NSFW ratio, defined as the proportion of all images flagged as NSFW and assigned any browsing level other than 1 (which is SFW), increased from 56% in October 2023 to 73% in April 2024. Additionally, Wei et al. (2024) demonstrated that NSFW images receive more social engagement on CivitAI than SFW images. Our current study extends the dataset to include model metadata and covers a full two-year period, from January 2023 to December 2024. With this extended dataset, we confirm a continuation of the trend, with the NSFW ratio increasing from 41% in January 2023 to 80% by the end of December 2024. This trend is visualized in Figure 4, which presents a stacked bar diagram showing the color-coded proportions of over 40 million images according to their assigned NSFW browsing level, posted to CivitAI between January 2023 and December 2024. A more detailed account of monthly trends in image generation and NSFW classification is provided in Table S1 in the Supplementary Material.
@@
Figure 4: Inferred age and gender predictions using MiVOLO on a 0.1% sample of temporally representative images from the 40 million image dataset.
@@ 2.2.2 Demographic Patterns in Generated Images
To analyze and quantify the demographic distribution of human subjects depicted in generated images shared via CivitAI, we employed the MiVOLO (Multi-Input VOLO) model introduced by Kuprashevich and Tolstykh (2023). This is a transformer-based system designed to predict age and gender from images, even when faces are partially or fully occluded, while excluding images without human subjects. The results are shown in Figure 4, which shows the distribution of estimated gender and ages and the corresponding NSFW browsing levels.
@@
The gender and age estimation was performed on a sampled subset of the original 40-million dataset. The subset consists of 40,636 images, representing approximately 0.1% of the full dataset. It was created by sampling, ensuring a temporally balanced and distributionally representative sample. Using the MiVOLO classifier, we identified 29,073 human subjects in 40,636 images, discarding images with no human subjects or a low confidence level for gender classification. The results of gender classification indicate a stark over-representation of female subjects, with an average female-to-male ratio of 6.24 to 1 across the dataset. NSFW co-occurrence rates were also notably higher for female subjects, with 85.44% of images depicting women with a browsing level other than 1, compared to male subjects, where the NSFW incidence was 73.66%. The average age for female subjects was 23.67 years, with a standard deviation of 5.1, while male subjects had a higher average age of 31.16 years with a standard deviation of 9.09 years. This indicates a narrower age range and a strong emphasis on sexualized portrayals of younger women, in contrast to a broader age distribution and more varied depictions of male individuals. A more detailed overview of these results, including a monthly breakdown of estimated gender and age distributions in sampled images, proportions across NSFW browsing levels, female-to-male ratios, and mean age statistics for each group is given in Table S2 in the Supplementary Material.
@@
Several limitations should be noted in this analysis: CivitAI assigns browsing levels using Amazon Rekognition, unspecified open-source NSFW classifiers, and manual reviews. NSFW classifiers disproportionately misclassify images of female subjects as explicit compared to male subjects (Leu et al., 2024), potentially skewing NSFW label distributions towards higher-sensitivity categories for females. Additionally, MiVOLO’s binary gender classification relies on visually inferred traits, ignoring non-binary and gender-diverse identities, and may produce uneven error rates, especially for darker-skinned individuals (Buolamwini and Gebru, 2018). Age estimation models also introduce biases; for instance, digital ageism can distort the representation of older adults, particularly older women (Chu et al., 2022).
@@ 2.3 Model-Level Analysis: Themes and Biases
While generated images often attract the most attention, it is important to recognize that platforms like CivitAI primarily function as model-sharing hubs. The distribution and adaptation of models can reinforce ethically concerning trends, such as the hypersexualization and objectification of women, often more systematically than individual images. Accordingly, we shifted focus from shared images to the generative models behind them. Of the 232,164 assets published on CivitAI between November 2022 and January 2025, only 4.28% were standalone checkpoint models, while 94.04% were model adapters, which require a compatible checkpoint to generate images. A breakdown of asset types appears in Figure S12 in the Supplementary Material. To examine the thematic focus of these models, we began with an analysis of user-generated promotional tags. We then evaluated gender bias reproduction by running inference tests on the ten most popular checkpoint models. Next, we examined thirty popular model adapters, such as LoRA models, assessing their function in augmenting the visual repertoire of base models. Within this group, we identified a specific subset, referred to as deepfake adapters, which were designed to replicate particular individuals and made up about 15% of all adapters.
@@ 2.3.1 Thematic Overview of Models shared on CivitAI
To understand dominant thematic orientations, we collected promotional tags added by model creators on CivitAI. These tags categorize and describe thematic, aesthetic, and functional attributes of models and adapters during the publishing process. We analyzed their frequency and co-occurrence across the platform. Figure 5 (interactive version available [^4] An interactive version of this graph is available here: https://pm-paper-viz.github.io/visualizations/figure_5.html) presents a co-occurrence network of the 300 most common tags. Each node represents a tag, with size indicating frequency. Edges indicate co-occurrence, with thicker, darker lines representing stronger connections. To aid visual interpretation, the size of nodes was normalized, and the top three most frequent nodes indicate their absolute occurrence counts. Anime aesthetics overall appeared to be a prevalent theme on the platform, with the tag “anime” appearing in roughly 30% of models. Additionally, tags denoting gender were highly prevalent. The tag “woman” appeared in 29.2% of shared models, “girls” in 15.7%, and “female” in 11.2%. Taken together, these figures indicate a pronounced emphasis on women and girls as depicted subjects across all assets exchanged on the platform. These tags were also commonly associated with tags such as “sexy”, “pose”, certain hairstyles and attire, and visual characteristics such as portraiture. The most active cluster revolved around character and anime and was connected to franchises like Dragon Ball, One Piece, and Gundam, and descriptors like TV show, visual novel, and game character. This cluster reflected the strong influence of fan culture and pop media. Among the various thematic patterns observed, the prominence of gendered tags stood out as one of the most striking results, highlighting deeper structural issues within the platform’s content ecosystem. Tags such as “woman” and “girl” not only dominated quantitatively but were also frequently linked to sexualized terms.
@@ 2.3.2 Bias in Model Checkpoints
To visualize what embedded discrimination in popular diffusion models means in terms of visual output, we performed an experiment on the top ten most popular checkpoints ranked by download count. Checkpoints are saved versions of a pretrained model, containing all necessary parameters to perform inference tasks without additional training and serving as the foundational assets for image synthesis, whereas adapters such as LoRA function as extensions merged into checkpoints to tweak their generative capabilities. To assess gender bias in popular checkpoint models shared on CivitAI, we performed a simple inference experiment based on comparing outputs of two different neutral text prompts. We selected the ten most popular checkpoint models according to download count available on the platform as of January 19th, 2025. For instance, Realistic Vision, the most downloaded asset at the time, had been downloaded over 1.5 million times, illustrating the scale of model adoption and its potential impact on generated content. We prompted each checkpoint model separately with the single-word prompts “woman” and “man”. Prompt embeddings were generated using the standard CLIP text encoder expected by each checkpoint. All images were generated using consistent parameters: CFG scale 6.5, Euler sampler (Qi et al., 2024), and 15 inference steps. To ensure comparability across outputs, we initialized the generation process with the same random noise seed for each model.
@@
The selected checkpoints and their inference results are visualized in Figure 6. For comparison, we also included results from Stable Diffusion 1.5 and Stable Diffusion XL 1.0, since most of the selected checkpoints, (aside from Pony Diffusion), are derivatives of these models. The generated outputs across all models revealed a stark difference in gender representation: women were predominantly depicted nude, while men were frequently portrayed in professional attire.
While the baseline models, Stable Diffusion 1.5 and Stable Diffusion XL, still exhibit noticeable gender biases, these biases are even more pronounced in the model derivatives popular on CivitAI. It must be noted that our selection of checkpoint models was based on platform popularity metrics without restricting for model generality or application domain. As a result, the sample includes both general-purpose models and specialized models (e.g., anime-stylized, NSFW-oriented). This choice reflects the actual diversity of models encountered by end-users on CivitAI and allows us to capture platform-wide trends in gender representation. However, it also introduces specific variations that may intensify gender biases. Taken together, these findings point toward embedded discrimination within the platform’s model ecosystem, where user preferences, model fine-tuning practices, and training data biases converge to systematically reproduce and amplify narrow and hypersexualized portrayals of women.
@@ 2.3.3 Most Popular Model Adapters
A brief examination of the promotional tags of the most popular models confirms the thematic orientation identified in our broader tag analysis, with frequent references to tags such as “anime”, “female” and “sexy”, suggesting a strong emphasis on gendered and sexualized representations. Indeed, a closer look at the 30 most downloaded adapters reveals a strong presence of adapters implying pornographic themes. Many of the promotional tags accompanying these models specifically emphasize female subjects or body parts, with three models focused solely on the depiction of vaginas. Disconcertingly, one of the most popular adapters, Nudify XL: Better Bodies, directly references the practice of non-consensual “nudification” where photos of dressed individuals are manipulated to appear naked. Surprisingly, even models focused on stylistic variation often include tags like “female”. While CivitAI does not disclose how tags influence visibility, it is likely that popular tags improve discoverability. This can create a feedback loop: widely used terms like “female”, “girl”, or “sexy” not only reflect user interest but also shape platform norms, as creators adopt them to boost exposure.
@@ 2.3.4 Deepfake Adapters and Targeted Subjects
One of the most ethically concerning applications of model personalization is the creation of model adapters intended to replicate real individuals. To better understand the scope and impact of this phenomenon, we focused on analyzing the subset of such adapters, investigating their prevalence, common use cases, and the demographic characteristics of the individuals they target. According to Deeptrace’s 2019 report, 96% of all deepfake content online was pornographic (Ajder et al., 2019), a trend that intensified by 2023 when a report by Home Security Heroes found that 98% of deepfake videos were pornographic and 99% of the victims were women (Home Security Heroes, 2023; Birrer and Just, 2024).
Although CivitAI prohibits users from publicly sharing explicit content involving real people, it has no control over how models are used once downloaded or what is shared on external platforms. This creates a significant blind spot in the enforcement of ethical guidelines, as deepfake models can still be deployed in harmful ways beyond the platform’s visibility. Given this regulatory gap and the continuing risk of abuse, we sought to understand who is being targeted by these models. To gain more insight into the individuals currently targeted by the practice of non-consensual deepfakes created from models shared on CivitAI, we annotated our dataset of deepfake adapters using a large language model, similarly to the approach pioneered by Wei et al. (2024). This allowed us to identify the most commonly associated professions and countries linked to those individuals.
@@
To visualize our results, we use sunburst charts, which represent nested categories as concentric rings, to illustrate how inferred professions are distributed across gender and country associations. Figure 8 includes two sunburst charts that visualize the distribution of inferred professions among individuals represented in the dataset. Figure 8(a) shows this distribution aggregated by gender. While Figure 8(b) breaks down the same professional categories by the most frequently associated countries.
Our results indicate that actors emerged as the most common profession across both genders and all countries, with the exception of Japan. Among women, adult performers were the second most frequent category, whereas public figures ranked second among men.
The majority of targeted individuals were associated with countries in the Global North, specifically the United States, the United Kingdom, and Japan. Cultural factors appear to influence who is targeted: for example, Japan showed a notably high presence of voice actors compared to other countries, likely reflecting the country’s strong anime and gaming industries. Sports professionals were also frequently identified, a pattern that may be attributed to the increased media exposure surrounding international events such as the 2024 Olympics.
This analysis offers insight into the cultural and professional context of deepfake victims but is limited by simplifying assumptions and methodological constraints. Language models may misclassify individuals or overlook multiple roles, and restricting profession predictions to three options while using only the first simplifies complex identities. DeepSeek can conflate distinct geographic regions. Nonetheless, the analysis provides a structured entry point for exploring the social dynamics behind deepfake creation.
The creation of deepfake explicit images is a clear abuse that highlights the risks of model personalization. Yet beyond these malicious uses lies a deeper, systemic problem: the spread of misogynistic and ethically troubling content through the very structures of model training. These issues often arise not from intent, but from seemingly neutral processes like image-text pairing. In the next section, we examine how such mechanisms embed gendered and sexualized assumptions into personalized models.